OpenStudy (anonymous):

How do I find the percent of discount to the nearest percent if my problem is... Hockey ticket: regular price, $25 sale price, $20

OpenStudy (anonymous):

I would assume it would be \[\frac{ saleprice }{ originalprice }\times 100\]

OpenStudy (anonymous):

Thank u

OpenStudy (anonymous):

anytime

OpenStudy (phi):

the percent discount will be 100 minus that value. for example, if you had something that normally sells for 100, and it is sold at a discount for 90, your formula will give you 90/100 = 90% the discount will be 10% (10% discount means take 10% off the original price.) you could find the discount percent by using (original - sales)/ original and then changing that fraction to a decimal, and then a percent.

OpenStudy (anonymous):

Thank u both :)

OpenStudy (phi):

what did you get for an answer?